OpenStudy (anonymous):

x^2+y^2-3x-4y+4=0 How is this broken down for a circle

OpenStudy (anonymous):

first seperate your variables from each other so put all your x's and y's next to each other and then subtract 4

OpenStudy (anonymous):

\[(x^2-3x)+(y^2-4y)=-4\]

OpenStudy (anonymous):

then complete the squares for x and y

OpenStudy (anonymous):

do you know how to do this

OpenStudy (anonymous):

I am not real sure how to do this.

OpenStudy (anonymous):

Completing the Square method will be used here..

OpenStudy (anonymous):

completing the square happens when you have an equation such as \[x^2+bx=0\] and you want to make it into a perfect square. to do this you take the coefficient b and divide it by two. after dividing it, square it and add it to both sides of the equation. the end product looks like this \[x^2+bx+\frac{b^2}{2}=\frac{b^2}{2}\]

OpenStudy (anonymous):

there is a mistake in the end... the form should be this \[x^2+bx+(\frac{b}{2})^2=(\frac{b}{2})^2\]

OpenStudy (anonymous):

the left side can be factored into a perfect square \[(x+\frac{b}{2})^2\]=(\frac{b}{2})^2\]

OpenStudy (anonymous):

so if you do that with both x and y what do you gte

OpenStudy (anonymous):

confused

OpenStudy (anonymous):

Thanks for the help!!!